A gas, initially at 292K, enclosed in a bulb of volume 15.1mL at a pressure of 90.2kPa, was first expanded into a reaction vessel so that the total volume was 182.2mL and then heated to 428K. What is the final pressure in the vessel?
1
Expert's answer
2020-09-29T07:00:56-0400

pV=nRT. n= pV/RT

n= 90.2 x 15.1/8.314 x 292 = 0.561.

p= nRT/V = 0.561 x 8.314 x 428 / 182.2 = 10.957 kPa.